(BASIC OPERATIONS WITH FRACTIONS)
How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 3 1/2 yards long, and 5 pieces, each of 2 1/4 yards long are removed?

Ans I trieed to do is * the 24 yard to 3 pieces but did not get the answer


ANS OF THIS PROBLEM IS 2 1/4
:
(BASIC OPERATIONS WITH FRACTIONS)
How many yards of material from a 24-yard length of cloth remain after 3 pieces, each 3 1/2 yards long, and 5 pieces, each of 2 1/4 yards long are removed?

Ans I trieed to do is * the 24 yard to 3 pieces but did not get the answer


ANS OF THIS PROBLEM IS 2 1/4

Answer by checkley77(3412)

24=3*3.5+5*2.25+X
24=10.5+11.25+X
24=21.75+X
X=24-21.75
X=2.25 OR 2 1/4 YARDS REMAIN.